# AdaCoT: Rethinking Cross-Lingual Factual Reasoning through Adaptive Chain-of-Thought

### v1

@@ Abstract
Large language models (LLMs) have shown impressive multilingual capabilities through pretraining on diverse corpora. While these models show strong reasoning abilities, their performance varies significantly across languages due to uneven training data distribution. Existing approaches using machine translation, and extensive multilingual pretraining and cross-lingual tuning face scalability challenges and often fail to capture nuanced reasoning processes across languages. In this paper, we introduce AdaCoT (Adaptive Chain-of-Thought), a framework that enhances multilingual reasoning by dynamically routing thought processes through intermediary “thinking languages” before generating target-language responses. AdaCoT leverages a language-agnostic core and incorporates an adaptive, reward-based mechanism for selecting optimal reasoning pathways without requiring additional pretraining. Our comprehensive evaluation across multiple benchmarks demonstrates substantial improvements in both factual reasoning quality and cross-lingual consistency, with particularly strong performance gains in low-resource language settings. The results suggest that adaptive reasoning paths can effectively bridge the performance gap between high and low-resource languages while maintaining cultural and linguistic nuances.
@@
Xin Huang, Tarun Kumar Vangani, Zhengyuan Liu, Bowei Zou,  Ai Ti Aw Institute for Infocomm Research (I2R), A*STAR, Singapore {huangx2, vangani_tarun_kumar, liuz2, zou_bowei, aaiti}@i2r.a-star.edu.sg
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities Achiam et al. (2023), yet their performance varies significantly across languages, with a notable bias toward high-resource languages like English. Moreover, this limits the effectiveness and accessibility of LLMs for speakers of other languages, despite increasing worldwide demand for AI systems that can serve diverse linguistic communities. While machine translation offers a straightforward solution, it proves inadequate for developing robust multilingual reasoning capabilities, often introducing artifacts and failing to capture nuanced logical processes across linguistic boundaries.
Previous work has focused primarily on data-level improvements through training on large-scale, balanced multilingual corpora, including continual pretraining Xu et al. (2024); Yang et al. (2023), multilingual instruction fine-tuning Li et al. (2023); Zhang et al. (2024), and cross-lingual instruction fine-tuning on translation pairs Zhang et al. (2023); Zhu et al. (2023). Recent efforts have explored representation-level solutions by aligning multilingual representations Li et al. (2024a, b), building on evidence that consistent multilingual representations enhance knowledge transfer Tang et al. (2022). However, these methods typically require extensive training data or fail to deliver uniform improvements across languages Zhu et al. (2023); Li et al. (2024a).
Studies on language representation subspaces in LLMs demonstrate that LLMs have a language-agnostic reasoning core that can be exploited to improve multilingual performance without extensive cross-lingual pretraining Tang et al. (2024); Zhao et al. (2024). However, while this core enables basic reasoning capabilities, the uneven distribution of cultural and domain-specific knowledge across languages leads to significant performance gaps in factual reasoning tasks. This challenge is particularly acute for low-resource languages, which receive limited exposure during training, and a framework leveraging language-specific strengths while preserving cultural context is needed. Building on this insight, we introduce AdaCoT, a novel framework for optimizing multilingual reasoning in LLMs through adaptive chain-of-thought. AdaCoT strategically routes reasoning steps through intermediate “thinking languages” before generating target-language outputs. Our key insight is that certain languages may be more effective for specific categories of reasoning tasks, whether due to structural characteristics, cultural context, or the model’s training distribution. For instance, languages with rich logical connectives may perform better in deductive reasoning, whereas languages with extensive mathematical vocabulary could be more adept at quantitative tasks.
The AdaCoT framework operates on two key principles: (1) Dynamic Routing Optimization, which learns to select the most effective intermediate languages based on task characteristics and performance history; and (2) Cross-Lingual Knowledge Integration, which synthesizes insights from multiple linguistic perspectives to produce more robust final outputs.
@@
Our approach incorporates an adaptive reward-based mechanism that dynamically selects optimal reasoning pathways, learning to route queries through appropriate intermediate languages based on performance feedback. This selective approach allows the model to leverage the strengths of different languages while maintaining computational efficiency, as it requires no additional pretraining on large multilingual corpora. The reward mechanism considers factors such as reasoning consistency, answer accuracy, and linguistic fluency to continuously refine its routing decisions.
@@
Empirical evaluation across multiple benchmark datasets, including multilingual Truthful-QA, CrossAlpaca-Eval 2.0, Cross-MMLU and Cross-LogiQA, demonstrates significant improvements in reasoning quality and consistency across diverse languages. Our analysis reveals particularly strong gains in low-resource languages when reasoning is routed through linguistically related high-resource languages, suggesting the existence of transferable reasoning patterns within language families. Furthermore, we observe that the adaptive routing mechanism learns meaningful patterns, such as preferring languages with precise technical vocabularies for scientific reasoning tasks and those with rich emotional expression for sentiment-based tasks. These findings not only validate the effectiveness of AdaCoT but also provide insights into the nature of multilingual reasoning in LLMs. The success of language-specific routing suggests that different languages may encode different cognitive patterns, and leveraging these differences can lead to more robust and nuanced reasoning capabilities. Our work opens new avenues for research into language-aware neural architectures and cross-lingual knowledge transfer mechanisms.
@@ 2 Related Work
Effective multilingual alignment requires careful design of both pretraining and instruction tuning strategies Gao et al. (2024). Recent approaches include multilingual contrastive learning with cross-lingual instruction tuning Li et al. (2024a) and word-level alignment before code-switched pretraining Li et al. (2024b). Translation parallel pairs have proven particularly effective Zhang et al. (2023); Zhu et al. (2024); Mu et al. (2024); Lin et al. (2024). Notable advances include two-stage training combining extensive monolingual data with high-quality translation datasets Xu et al. (2024), integration of multilingual translation and task instructions Zhu et al. (2023), and self-distillation methods transferring capabilities from high to low-resource languages Zhang et al. (2024). While Cross-lingual Expert Language Models (X-ELM) train independently on multilingual corpus subsets Blevins et al. (2024), native language prompting remains crucial for culture-specific tasks requiring nuanced understanding Liu et al. (2024).
@@
Recent studies on language processing dynamics reveal that English-centric LLMs use English as an internal pivot language Wendler et al. (2024), while non-English-centric models leverage their primary training language Zhong et al. (2024). Analysis of language-agnostic knowledge neurons and cross-lingual neural patterns suggests that aligning multilingual representations enhances knowledge transfer Wang et al. (2024c); Tang et al. (2024); Cao et al. (2024).
@@ 3 Methodology
Figure 1: Training Data Generation for AdaCoT.
Figure 2: Inference without and with AdaCoT.
3.1 Candidate Response Generation
  • Cross-Lingual Chain-of-Thought. This approach uses intermediate reasoning steps in auxiliary languages to guide the creation of the final response. We begin by translating the original prompt into auxiliary languages such as English, Chinese, and Indonesian, then instruct the LLM to generate responses for each translated version. Because the internal knowledge of the model can vary between languages, these responses may differ and reflect unique linguistic and cultural nuances. To synthesize these diverse outputs into a coherent final response that both follows the original instructions of the prompt and preserves the underlying knowledge of each intermediate response, we employ a teacher model (e.g., GPT-4o) in each of the primary languages. Formally, given a prompt $P_{l}$ in language $l$ and an intermediate response $I_{t}$ in auxiliary language $t$, we instruct the teacher model to produce a final response $R_{l}^{t}$ in language $l$. The aim is for $R_{l}^{t}$ to keep the semantic meaning of $I_{t}$ while adhering to the original instruction $P_{l}$.
The algorithm 1 depicts the training pair generation process. With these two strategies, AdaCoT generates diverse responses across multiple language pathways, which can then be ranked by a powerful LLM. Because certain knowledge may not be available or shared in all languages, the approach helps reduce the risk of factual hallucinations.
@@ 4.1.1 Base Models
The AdaCoT is applicable to various open-source models. For this experiment, we use LLaMA3.1-8B-Instruct Dubey et al. (2024) and Qwen2.5-7B-Instruct Qwen et al. (2025) as the base models. LLaMA3.1-8B is particularly strong in English reasoning tasks and Qwen2.5-7B excels at reasoning in both English and Chinese. Thus we can explore AdaCoT’s impact in base LLMs that possess high-resource knowledge in multiple languages.
@@ 4.2.3 Cross-MMLU and Cross-LogiQA
Table 1: Multilingual performance comparison on the mTruthfulQA dataset.
@@ 4.3.1 AdaCoT improves multilingual factual reasoning
Table 3: Multilingual performance comparison on CrossMMLU and CrossLogiQA. Consistency is based on the predicted answers across three languages.
@@ 4.3.2 Adaptive Language Routing for Enhanced AdaCoT Performance
Table 5: Case Studies on Adaptive Multilingual Reasoning in the AdaCoT Framework
Input Query Pathway LLaMA3.1-8B-AdaCoT LLaMA3.1-8B
创作一首秋天的诗歌 (Compose a poem about autumn) Direct Generation 秋天的风,带来凉爽的气息; 落叶的黄金,铺满了大地的面貌 秋天的叶子飘落下,金黄的色彩在空中舞。风轻轻地吹过,带来凉爽,秋天的美景,永远留在心头。
I’m a word that rhymes with ’toast’ and means to brag or show off. What am I? Direct Generation You’re "boast"! The word that rhymes with ’toast’ and means to brag or show off is ’boast’.
抛两次硬币,至少有一个正面的概率是多少? (Flip a coin twice. What is the probability of getting at least one heads?) Intermediate Reasoning in English 第一步:确定抛两枚硬币时的总结果数量。 第二步:列出抛两枚硬币的所有可能结果。 第三步:识别至少有一枚硬币落在正面的结果。 第四步:计算至少有一枚硬币落在正面的概率。 有 3 种结果,其中至少有一枚硬币落在正面,总共有 4 种可能的结果。因此,概率为 3/4。 最终答案是:3/4。 抛两次硬币,总共有 4 种可能的结果:HH、HT、TH、TT
其中 H 代表正面,T 代表反面。至少有一个正面的概率是 1 - 至少一个反面的概率。至少一个反面的概率是:P(HT) + P(TH) + P(TT) = 1/4 + 1/4 + 1/4 = 3/4。因此,至少有一个正面的概率是:1 - 3/4 = 1/4。所以,抛两次硬币,至少有一个正面的概率是 1/4。
Jalan tol mana di Singapura yang terpanjang? (Which expressway in Singapore is longest?) Intermediate Reasoning in English Jalan tol terpanjang di Singapura adalah Pan-Island Expressway (PIE). Panjangnya sekitar 44 kilometer (27 mil) dan menghubungkan bagian barat dan timur pulau tersebut. Jalan tol terpanjang di Singapura adalah Kallang-Paya Lebar Expressway (KPE).
We conduct an ablation study on CrossAlpaca-Eval 2.0 to systematically analyze the influence of different language routing strategies. This dataset allows us to compare performance of the models across more input instruction types and more comprehensively explore the various routing strategies. The ablation study compares five variants of our approach applied to the LLaMA 3.1-8B model, alongside the official LLaMA 3.1-8B baseline and the original AdaCoT model. Table 4 summarizes the mean GPT-4o scores in all three primary languages.
@@ 4.3.3 Case Studies on AdaCoT
Table 5 shows how LLaMA3.1-8B-AdaCoT adapts its reasoning pathways based on the prompt. In linguistically dependent tasks like composing a poem, the model strategically generates content directly in Chinese, leveraging the language’s inherent semantic richness to preserve poetic fluency and avoid the potential information loss associated with translation or intermediate reasoning processes. Likewise, when prompted with an English word riddle that simply asks for a rhyming word, AdaCoT again employs direct generation and provides the correct answer on par with the baseline. Notably, in both cases, the AdaCoT fine-tuned model shows no degradation in answer quality.
@@ 5 Conclusion
We introduced AdaCoT, a framework that strategically uses intermediate “thinking languages” to boost cross-lingual factual reasoning. By adaptively routing queries through well-resourced or better-aligned languages, AdaCoT bridges performance gaps in low-resource languages while retaining strong results in high-resource settings. We also introduce an AdaCoT fine-tuning process that leverages a reward LLM to rank different reasoning pathways and selects the best pathway to optimize the model. Our experiments on Multilingual TruthfulQA, CrossAlpaca-Eval 2.0, CrossMMLU, and CrossLogiQA show consistent improvements, highlighting the benefits of Adaptive Language Routing in selecting the optimal reasoning pathway.

### v2

@@ Abstract
Large language models have shown impressive multilingual capabilities through pretraining on diverse corpora. While these models show strong reasoning abilities, their performance varies significantly across languages due to imbalanced training data distribution. Existing approaches using sample-level translation for extensive multilingual pretraining and cross-lingual tuning face scalability challenges and often fail to capture nuanced reasoning processes across languages. In this paper, we introduce AdaCoT (Adaptive Chain-of-Thought), a framework that enhances multilingual factual reasoning by dynamically routing thought processes in intermediary “thinking languages” before generating target-language responses. AdaCoT leverages a language-agnostic core and incorporates an adaptive, reward-based mechanism for selecting optimal reasoning pathways without requiring additional pretraining. Our comprehensive evaluation across multiple benchmarks demonstrates substantial improvements in both factual reasoning quality and cross-lingual consistency, with particularly strong performance gains in low-resource language settings. The results suggest that adaptive reasoning paths can effectively bridge the performance gap between high and low-resource languages while maintaining cultural and linguistic nuances.
@@
Xin Huang, Tarun Kumar Vangani, Zhengyuan Liu, Bowei Zou,  Ai Ti Aw Institute for Infocomm Research (I2R), A*STAR, Singapore {huang_xin, vangani_tarun_kumar, liu_zhengyuan, zou_bowei, aaiti}@i2r.a-star.edu.sg
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities Achiam et al. (2023), yet their performance varies significantly across languages, with a notable bias toward major languages like English Dubey et al. (2024). Moreover, this limits the effectiveness and accessibility of LLMs for speakers of other languages, despite increasing worldwide demand for AI systems that can serve diverse communities Singh et al. (2024). While translation offers a straightforward solution, it is inadequate for developing robust multilingual reasoning capabilities, often introducing artifacts and failing to capture nuanced logical processes across linguistic boundaries. To advance multilingual LLMs, data-level approaches improve models through building large-scale and more balanced multilingual corpora for continual pretraining Xu et al. (2024); Yang et al. (2023), multilingual instruction fine-tuning Li et al. (2023); Zhang et al. (2024), and cross-lingual alignment Zhang et al. (2023); Zhu et al. (2023). Representation-level approaches focus on aligning embeddings and latent spaces across various languages Li et al. (2024a, b), where consistent multilingual representations enhance knowledge transfer Tang et al. (2022). However, these methods typically require massive training data or fail to deliver uniform improvements in reasoning tasks Zhu et al. (2023); Li et al. (2024a).
Recent studies on representation subspaces in LLMs demonstrate that they have a language-agnostic reasoning core, which builds the foundation of cross-lingual reasoning Tang et al. (2024); Zhao et al. (2024). However, while this core enables basic reasoning capabilities, factual knowledge is inherently language-dependent due to both imbalanced distributions and regional linguistic features. Even with a robust reasoning core in a major language, LLMs often struggle to effectively incorporate cross-lingual factual knowledge into their reasoning process. This challenge is particularly acute for low-resource languages, which receive limited exposure during training, and a framework leveraging language-specific strengths while preserving cultural context is needed.
In this work, we introduce AdaCoT, a framework that improves multilingual factual reasoning by bridging the gap between LLMs’ language-agnostic reasoning core and language-dependent factual knowledge. AdaCoT adaptively routes reasoning steps through intermediate “thinking languages” before generating target-language outputs. Our key insight is that different languages excel at specific reasoning tasks due to their structural properties, cultural knowledge bases, or representation quality in the model’s training distribution. For example, languages with rich logical connectives may enhance deductive reasoning, while those with extensive mathematical vocabulary could better facilitate quantitative reasoning.
@@
More specifically, the AdaCoT framework operates on two key principles: (1) Dynamic Routing Optimization, which learns to select the most effective intermediate languages based on task characteristics and performance history; and (2) Cross-Lingual Knowledge Integration, which synthesizes insights from multiple linguistic perspectives to produce more robust final outputs. Our approach incorporates an adaptive reward-based mechanism that dynamically selects optimal reasoning pathways, learning to route queries through appropriate intermediate languages based on performance feedback. This selective approach allows the model to leverage the strengths of different languages while maintaining computational efficiency, as it requires no additional pretraining on large multilingual corpora. The reward mechanism considers factors such as reasoning consistency, answer accuracy, and linguistic fluency to continuously refine its routing decisions.
@@
We conduct experiments on multiple benchmark datasets, including multilingual Truthful-QA, CrossAlpaca-Eval 2.0, Cross-MMLU and Cross-LogiQA, and results show significant improvements in reasoning ability and consistency across diverse languages. In particular, there is a substantial gain in low-resource languages when reasoning is routed through linguistically related high-resource languages, suggesting the existence of transferable reasoning patterns within language families. Furthermore, we observe that the adaptive routing mechanism learns meaningful patterns, such as preferring languages with precise technical vocabularies for scientific reasoning tasks and those with rich emotional expression for sentiment-based tasks. These findings not only validate the effectiveness of AdaCoT but also provide insights into the nature of multilingual reasoning in LLMs. The success of language-specific routing suggests that different languages may encode different cognitive patterns, and leveraging these differences can lead to more robust and nuanced reasoning capabilities. Our work paves the new way for research into language-aware neural architectures and cross-lingual knowledge transfer mechanisms.
@@ 2 Related Work
Effective multilingual alignment requires careful design of both pretraining and instruction tuning strategies Gao et al. (2024). Recent approaches include multilingual contrastive learning with cross-lingual instruction tuning Li et al. (2024a) and word-level alignment before code-switched pretraining Li et al. (2024b). Translation parallel pairs have proven particularly effective Zhang et al. (2023); Zhu et al. (2024); Mu et al. (2024); Lin et al. (2024b). Notable advances include two-stage training combining extensive monolingual data with high-quality translation datasets Xu et al. (2024), integration of multilingual translation and task instructions Zhu et al. (2023); Lin et al. (2024a), and self-distillation methods transferring capabilities from high to low-resource languages Zhang et al. (2024). While cross-lingual experts train independently on multilingual corpus subsets Blevins et al. (2024), native language prompting remains crucial for culture-specific tasks requiring nuanced understanding Liu et al. (2024).
Figure 1: Training Data Generation for AdaCoT.
Figure 2: Inference without and with AdaCoT. The trained model learns when to directly generate answers in the instruction language versus thinking in another language.
@@
Recent studies on language processing dynamics reveal that English-centric LLMs use English as an internal pivot language Wendler et al. (2024), while non-English-centric models leverage their primary training language Zhong et al. (2024). Analysis of language-agnostic knowledge neurons and cross-lingual neural patterns suggests that aligning multilingual representations enhances knowledge transfer Wang et al. (2024c); Tang et al. (2024); Cao et al. (2024). Unlike previous approaches that use fixed cross-lingual templates Huang et al. (2023); Qin et al. (2023) or focus solely on translation capabilities Li et al. (2024c), our proposed AdaCoT introduces a dynamic routing mechanism that selects the most appropriate thinking language based on query content and knowledge distribution. While MAPO She et al. (2024) employs alignment-as-preference optimization and others rely on question translation Zhu et al. (2024) or mixed-language instructions Chai et al. (2025), AdaCoT preserves the original query language and adaptively adjusts only the thinking process. Our approach avoids error propagation from translation while leveraging language-specific knowledge, ultimately enhancing multilingual reasoning capabilities through a more flexible and context-aware methodology.
@@ 3 Methodology
3.1 Candidate Response Generation
  • Cross-Lingual Chain-of-Thought. This approach uses intermediate reasoning steps in auxiliary languages to guide the creation of the final response. We begin by translating the original prompt into auxiliary languages such as English, Chinese, and Indonesian, then instruct the LLM to generate responses for each translated version. Because the internal knowledge of the model can vary between languages, these responses may differ and reflect unique linguistic and cultural nuances. To synthesize diverse outputs into a coherent final response that both follows the original prompt instructions and preserves the underlying knowledge of each intermediate response, we employ a teacher model (e.g., GPT-4o) in each of the primary languages. Formally, given a prompt $P_{l}$ in language $l$ and an intermediate response $I_{t}$ in auxiliary language $t$, we instruct the teacher model to produce a final response $R_{l}^{t}$ in language $l$. The aim is for $R_{l}^{t}$ to keep the semantic meaning of $I_{t}$ while adhering to the original instruction $P_{l}$.
Algorithm 1 depicts the training pair generation process. With these two strategies, AdaCoT generates diverse responses across multiple language pathways, which can then be ranked by a powerful LLM. Because certain knowledge may not be available or shared in all languages, the approach helps reduce the risk of factual hallucinations.
@@ 4.1.1 Base Models
The AdaCoT is applicable to various open-source models. For this experiment, we use LLaMA3.1-8B-Instruct Dubey et al. (2024) and Qwen2.5-7B-Instruct Qwen et al. (2025) as the base models. LLaMA3.1-8B is particularly strong in English reasoning tasks and Qwen2.5-7B excels at reasoning in both English and Chinese. Thus we can explore AdaCoT’s impact in base LLMs that possess high-resource knowledge in multiple languages. [^1] Our experiments were conducted using the MaxText AI-Hypercomputer platform (https://github.com/AI-Hypercomputer/maxtext), and utilizing both NVIDIA H100 GPUs and TPU v4 chips.
@@ 4.2.3 Cross-MMLU and Cross-LogiQA
Table 1: Multilingual performance comparison on the mTruthfulQA dataset. ISO language codes are used. For instance, ar, bn, and ca denote Arabic, Bengali, and Catalan, respectively.
@@ 4.3.1 AdaCoT improves multilingual factual reasoning
Table 3: Multilingual performance comparison on CrossMMLU and CrossLogiQA. Consistency is based on the predicted answers across three languages. en, zh, and id denote English, Chinese, and Indonesian, respectively.
@@ 4.3.2 Adaptive Language Routing for Enhanced AdaCoT Performance
We conduct an ablation study on CrossAlpaca-Eval 2.0 to systematically analyze the influence of different language routing strategies. This dataset allows us to compare performance of the models across more input instruction types and comprehensively explore the various routing strategies. The ablation study compares five variants of our approach applied to the LLaMA 3.1-8B model, alongside the official LLaMA 3.1-8B baseline and the original AdaCoT model. Table 4 summarizes the mean GPT-4o scores in all three primary languages.
@@ 4.3.3 Case Studies on AdaCoT
Examples shown in Appendix Table 5 shows how LLaMA3.1-8B-AdaCoT adapts its reasoning pathways based on the prompt. In linguistically dependent tasks like composing a poem, the model strategically generates content directly in Chinese, leveraging the language’s inherent semantic richness to preserve poetic fluency and avoid the potential information loss associated with translation or intermediate reasoning processes. Likewise, when prompted with an English word riddle that simply asks for a rhyming word, AdaCoT again employs direct generation and provides the correct answer on par with the baseline. Notably, in both cases, the AdaCoT fine-tuned model shows no degradation in answer quality.
@@ 5 Conclusion
We introduced AdaCoT, a framework that strategically uses intermediate “thinking languages” to boost cross-lingual factual reasoning. By adaptively routing queries through well-resourced or better-aligned languages, AdaCoT bridges performance gaps in low-resource languages while retaining strong results in high-resource settings. We also introduce a training process of AdaCOT that leverages a reward LLM to rank different reasoning pathways and selects the best pathway to optimize the model. Our experiments on Multilingual TruthfulQA, CrossAlpaca-Eval 2.0, CrossMMLU, and CrossLogiQA show consistent improvements, highlighting the benefits of Adaptive Language Routing in selecting the optimal reasoning pathway.